Need help on large speed bump. How to calculate standard pounds per unit and direct materials quantity variance?Last month, Banner Corporation purchased and used the same quantity of material in producing its product, speed bumps for tra
Answer #1
Given, SR/pound = $ 2.50
No. of units produced = 600 units
Actual quantity purchased and used per unit = 18
Actual price per material pound = $ 2.80
Direct material price variance = 3,240 (U)
Total Direct material variance = 1,740 (U)
Material Quantity varaince = Total Direct material variance - Direct material price variance
            = -1,740 + 3,240
            = 1,500 (F)
SQ * SR - AQ *SR = 1,500
SQ *2.5 = (600*18)*2.5 = 1,500
SQ = 11,400
Standard pounds / unit = 11,400/600 =19
Material Quantity variance = 1,500 (F)
Note :
SQ Standard quantity
SR Standard rate
AQ Actual Quantity